# The Very Faint X-ray Transient 4XMM J174610.7–290020 at the Galactic center

### v1

@@ 6 Discussion
4XMM J174610.7–290020 underwent two distinct outbursts in the last two years. Assuming that the source is located at the GC, the fluxes reported in Sect. 3 and 4 translates into a luminosity of $L_{2-10}\simeq(1-12)\times 10^{34}$ erg s-1 and $L_{2-10}\simeq(6-9)\times 10^{34}$ erg s-1 for the 2024 and 2025 outburst respectively. These two outbursts are the brightest ever recorded for this source; therefore, the transient belongs to the VFXT class. The same source was detected in the past at lower fluxes by Chandra and XMM-Newton. The flux in the previous detections is in the range $F_{2-10}\simeq(0.2-4)\times 10^{-13}$ erg cm-2 s-1 (see Fig. 7), corresponding to a luminosity $L_{2-10}\simeq(0.2-3)\times 10^{33}$ erg s-1, assuming the source is located at the GC. Upper limits instead indicates a quiescence luminosity of $L_{2-10}\leq 2\times 10^{31}$ erg s-1. Our spectral analysis of the 2024 outburst is in agreement with the results of Yoshimoto et al. (2025). When fitted with the same model (apec or power law continuum), the parameters are consistent (i.e., $kT\sim 12$ keV and $\Gamma\sim 2$), showing that there is no sign of spectral evolution during the outburst.
@@
Establishing the nature of this source is not straightforward. The spectrum is characterized by a spectral slope of $\Gamma\sim 2-3$, which is common for VFXT that are NS-LMXBs (Degenaar & Wijnands, 2009; Wijnands et al., 2015). Our analysis does not answer whether the iron complex emission is due to a single reflection line or the sum of ionized Fe XXV and Fe XXVI lines. In any case, iron lines are unusual for VFXTs which are NS-LMXBs, with only a few cases being observed (van den Eijnden et al., 2018; Del Santo et al., 2007). For example, van den Eijnden et al. (2018) reported for the VFXT IGR J17062–6143 the presence of a reflection iron line in the spectrum, which could arise in the case of a UCXB or a magnetically truncated disk.
@@
The study of the potential type I burst, is a strong indication of the NS-LMXB nature of the source. Indeed, the flare observed in 2004 shows several key features typical of type I bursts: the light curve (a fast rise followed by an exponential decay), the spectral shape (a blackbody with a temperature of a few keV), and the spectral softening during the burst (Galloway et al., 2008; Degenaar et al., 2018). Nevertheless, the duration of a few hundred seconds is longer than typical type I bursts, and falls in the intermediate-duration bursts class. Those bursts are related to NS-LMXB accreting below the Eddington limit (i.e., $0.03\,\dot{M}_{\mathrm{Edd}}$), and present a recurrence time of months to years rather than the hours to days of regular bursts (Alizai et al., 2023).
@@
A coherent explanation of these features can be achieved assuming that this particular VFXT is an accretion disk corona source, as proposed by Yoshimoto et al. (2025). In this scenario the NS-LMXB is seen almost edge on, and the radiation from the source is blocked by the thick accretion disk. The only observed flux is the emission scattered by the accretion disk corona, while the iron lines are due to the ionized plasma. In this case, both the outburst and the type I burst are intrinsically brighter, but only the fainter, scattered emission is observed.
@@
Other physical scenarios can explain part of the characteristics of this source, but struggle to reproduce all of them coherently. For example, HMXBs are another class of sources that exhibit ionized iron lines due to wind-fed accretion. HMXBs with X-ray luminosities comparable to those observed during the outburst of the last two years are known (Lutovinov et al., 2013; Mandel et al., 2025). However, it remains challenging to explain the burst feature with a softening tail that occurred during the XMM-Newton observations of 2004.
@@
CVs also present a thermal spectrum with strong ionized iron lines. In particular, the observed X-ray luminosity can be reached by an intermediate polar (IP). IPs are magnetic CVs that have non-synchronized orbits. The white dwarf’s magnetic field strength is $B\sim(0.1-10)\times 10^{6}\,\text{G}$; between polars ($B\sim(10-300)\times 10^{6}\,\text{G}$), and non-magnetic CVs ($B\lesssim 10^{5}\,\text{G}$). However, IPs are usually persistent and more luminous ($L_{2-10}\sim\times 10^{33}$ erg s-1, Mukai, 2017; Xu et al., 2019), therefore they cannot explain the observed broad variability of about 3 orders of magnitude and a quiescent luminosity $2\times 10^{31}$ erg s-1. In principle, a white dwarf can experience such luminosity variations in the case of a nova [^8] A nova is a thermonuclear runaway of H-rich material on the surface of a white dwarf. (Mukai, 2017). However, even if the two events observed in 2024 and 2025 were novae, it remains difficult to explain the rapid flare that occurred in 2004.
@@
Regarding the search for a NIR counterpart, it is inconclusive, as low mass main-sequence stars are practically non detectable because of source confusion at faint magnitudes. Regarding the four NIR sources which lie within the positional uncertainty of 4XMM J174610.7–290020, using the extinction value of each source reported in Nogueras-Lara et al. (2021) and assuming a distance of 8.2 kpc, the absolute K-magnitude of the four sources is in the range between $M_{K}\simeq-2.7$ and $M_{K}\simeq-0.5$. These magnitudes are typical of giants (see Nogueras-Lara et al. (2021) for details), or type-B stars (Mandel et al., 2025). The X-ray analysis suggests that the source is a NS-LMXB, therefore the NIR counterpart is likely undetected.
@@
If the low luminosity observed during the 2024 and 2025 events results from the accretion disk corona scenario, then 4XMM J174610.7–290020 would be classified as a VFXT due to its inclination rather than to a low accretion rate. However, as Fig. 7 shows, previous detections at fluxes 1–3 orders of magnitude lower suggest that the source also experienced phases of reduced accretion. For example, during the 2004 type I burst, immediately preceding the flare, the source was nearly two orders of magnitude fainter than during the 2024 and 2025 outbursts. The low accretion state in that period may therefore explain the unusually long duration of that type I burst (Alizai et al., 2023).
@@ 7 Conclusions
We analyzed the two distinct outbursts of the Galactic center VFXT 4XMM J174610.7–290020 (also known as XRISM J174610.8–290021) that occurred in 2024 and 2025. These represent the brightest outbursts ever observed from this source, lasting approximately 50 and 5 days, respectively, with a typical X-ray luminosity in the 2–10 keV band of $L_{2-10}\simeq(1-12)\times 10^{34}$ erg s-1. In both events, the source exhibited spectra characterized by a strong iron complex. An accretion disk corona, as proposed by Yoshimoto et al. (2025), can account for the transient’s low luminosity and its spectral shape. We studied a short flare from the same source that occurred in 2004, which shows the characteristics of type I X-ray bursts, thereby confirming its NS-LMXB nature. Although the accretion disk corona scenario can account for the relatively low luminosity of the recent outbursts, the source has previously been observed at luminosities up to three orders of magnitude lower, indicating intrinsically fainter outbursts, and a broad variability in its accretion behavior.
@@ Acknowledgements.
This work is based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. This project acknowledges financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program HotMilk (grant agreement No. 865637), support from Bando per il Finanziamento della Ricerca Fondamentale 2022 dell’Istituto Nazionale di Astrofisica (INAF): GO Large program and from the Framework per l’Attrazione e il Rafforzamento delle Eccellenze (FARE) per la ricerca in Italia (R20L5S39T9). SM acknowledges financial support through the INAF grants “Magnetars” and “Toward Neutron Stars Unification”. SM acknowledges support from the grant NASA ADAP 80NSSC24K0666. CJ acknowledges the National Natural Science Foundation of China through grant 12473016, and the support by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550200). LS acknowledges funding from the grant entitled “Bando Ricerca Fondamentale INAF 2023”

### v2

@@ 6 Discussion
Swift J174610-290018 underwent two distinct outbursts in the last two years. Assuming that the source is located at the GC, the fluxes reported in Sects. 3 and 4 translates into a luminosity of $L_{2-10}\simeq(1-12)\times 10^{34}$ erg s-1 and $L_{2-10}\simeq(6-9)\times 10^{34}$ erg s-1 for the 2024 and 2025 outburst, respectively. These two outbursts are the brightest ever recorded for this source; therefore, the transient belongs to the VFXT class. The same source was detected in the past at lower fluxes by Chandra and XMM-Newton. The flux in the previous detections is in the range $F_{2-10}\simeq(0.2-4)\times 10^{-13}$ erg cm-2 s-1 (see Fig. 7), corresponding to a luminosity $L_{2-10}\simeq(0.2-3)\times 10^{33}$ erg s-1, assuming the source is located at the GC. Upper limits instead indicates a quiescence luminosity of $L_{2-10}\leq 2\times 10^{31}$ erg s-1. Our spectral analysis of the 2024 outburst is in agreement with the results of Yoshimoto2025. When fitted with the same model (apec or power law continuum), the parameters are consistent (i.e., $kT\sim 12$ keV and $\Gamma\sim 2$), showing that there is no sign of spectral evolution during the outburst.
@@
Establishing the nature of this source is not straightforward. The spectrum is characterized by a spectral slope of $\Gamma\sim 2-3$, which is common for VFXT that are NS-LMXBs (Degenaar2009; Wijnands2015). Our analysis does not answer whether the iron complex emission is due to a single reflection line or the sum of ionized Fe XXV and Fe XXVI lines. In any case, iron lines are unusual for VFXTs, which are NS-LMXBs, with only a few cases being observed (Eijnden2018; Delsanto2007). For example, Eijnden2018 reported for the VFXT IGR J17062–6143 the presence of a reflection iron line in the spectrum, which could arise in the case of an ultra-compact X-ray binary or a magnetically truncated disk.
@@
The study of the potential type I burst, is a strong indication of the NS-LMXB nature of the source. Indeed, the flare observed in 2004 shows several key features typical of type I bursts: the light curve (a fast rise followed by an exponential decay), the spectral shape (a blackbody with a temperature of a few keV), and the spectral softening during the burst (Galloway2008; Degenaar2018). Nevertheless, the duration of a few hundred seconds is longer than typical type I bursts, and falls in the intermediate-duration bursts class. Those bursts are related to NS-LMXB accreting below the Eddington limit (i.e., $0.03\,\dot{M}_{\mathrm{Edd}}$), and present a recurrence time of months to years rather than the hours to days of regular bursts (Alizai2023).
@@
A coherent explanation of these features can be achieved assuming that this particular VFXT is an accretion disk corona source, as proposed by Yoshimoto2025. In this scenario the NS-LMXB is seen almost edge on, and the radiation from the source is blocked by the thick accretion disk. The only observed flux is the emission scattered by the accretion disk corona, while the iron lines are due to the ionized plasma. In this case, both the outburst and the type I burst are intrinsically brighter, but only the fainter, scattered emission is observed. In principle, if the NS–LMXB system is viewed almost edge-on, as proposed in this scenario, eclipses should be periodically visible in the light curve. The lack of detected eclipses may be due to the limited exposure time of the observations. In addition, the accretion disk corona can be larger than the line of sight obscured by the donor passage. For instance, for a NS-LMXB, with NS of $1.4\,M_{\odot}$, and donor star of mass $0.5\,M_{\odot}$, the Roche lobes are respectively $R_{L1}\simeq 0.8\,R_{\odot}$ and $R_{L2}\simeq 0.5\,R_{\odot}$ (Eggleton1983). Assuming that the extension of the accretion disk corona is roughly the circularization radius ($\approx 0.8\,R_{L1}$), the donor star radius is less than it. Therefore, the passage of the donor star never entirely obscure the extended source. Alternatively, since the type I burst implies that the source is a NS-LMXB, the lack of eclipses or dips would imply that the system is not observed edge-on and, therefore, that the source and the type I burst are intrinsically faint. The iron complex excess can then be a reflection disk line.
@@
Other physical scenarios can explain part of the characteristics of this source, but struggle to reproduce all of them coherently. For example, HMXBs are another class of sources that exhibit ionized iron lines due to wind-fed accretion. HMXBs with X-ray luminosities comparable to those observed during the outburst of the last two years are known (Lutovinov2013; Mandel2025). However, it remains challenging to explain the burst feature with a softening tail that occurred during the XMM-Newton observations of 2004.
@@
CVs also present a thermal spectrum with strong ionized iron lines. In particular, the observed X-ray luminosity can be reached by an intermediate polar (IP). IPs are magnetic CVs that have unsynchronized orbits. The white dwarf’s magnetic field strength is $B\sim(0.1-10)\times 10^{6}\,\text{G}$; between polars ($B\sim(10-300)\times 10^{6}\,\text{G}$), and nonmagnetic CVs ($B\lesssim 10^{5}\,\text{G}$). However, IPs are usually persistent and more luminous ($L_{2-10}\sim\times 10^{33}$ erg s-1; Mukai2017; Xu2019); therefore, they cannot explain the observed broad variability of about 3 orders of magnitude and a quiescent luminosity of $2\times 10^{31}$ erg s-1. In principle, a white dwarf can experience such luminosity variations in the case of a nova [^14] A nova is a thermonuclear runaway of H-rich material on the surface of a white dwarf. (Mukai2017). This scenario has been recently proposed by Hua2025. However, even if the two events observed in 2024 and 2025 were indeed novae, the rapid flare (type I burst like) observed in 2004 remains difficult to explain, strongly suggesting that the source is a NS-LMXB.
@@
Regarding the search for a NIR counterpart, it is inconclusive, as low mass main-sequence stars are practically non detectable because of source confusion at faint magnitudes. Regarding the four NIR sources that lie within the positional uncertainty of Swift J174610-290018, using the extinction value of each source reported in Nogueras2021 and assuming a distance of 8.2 kpc, the absolute K-magnitude of the four sources is in the range between $M_{K}\simeq-2.7$ and $M_{K}\simeq-0.5$. These magnitudes are typical of giants (see Nogueras2021 for details), or type-B stars (Mandel2025). The X-ray analysis suggests that the source is a NS-LMXB; therefore, the NIR counterpart is likely undetected.
@@
If the low luminosity observed during the 2024 and 2025 events results from the accretion disk corona scenario, then Swift J174610-290018 would be classified as a VFXT due to its inclination rather than to a low accretion rate. However, as Fig. 7 shows, previous detections at fluxes 1–3 orders of magnitude lower suggest that the source also experienced phases of reduced accretion. For example, during the 2004 type I burst, immediately preceding the flare, the source was nearly two orders of magnitude fainter than during the 2024 and 2025 outbursts. The low accretion state in that period may therefore explain the unusually long duration of that type I burst (Alizai2023).
@@ 7 Conclusions
We analyzed the two distinct outbursts of the GC VFXT Swift J174610-290018 (also known as XRISM J174610.8-290021) that occurred in 2024 and 2025. These represent the brightest outbursts ever observed from this source; they lasted approximately 50 and 5 days, respectively, with a typical X-ray luminosity in the 2–10 keV band ($L_{2-10}$) of $\simeq(1-12)\times 10^{34}$ erg s-1. In both events, the source exhibited spectra characterized by a strong iron complex. An accretion disk corona, as proposed by Yoshimoto2025, can account for the transient’s low luminosity and its spectral shape. We studied a short flare from the same source that occurred in 2004, which shows the characteristics of type I X-ray bursts, thereby confirming its NS-LMXB nature. Although the accretion disk corona scenario can account for the relatively low luminosity of the recent outbursts, the source has previously been observed at luminosities up to three orders of magnitude lower, indicating intrinsically fainter outbursts and a broad variability in its accretion behavior.
@@ Acknowledgements.
We thank the anonymous referee for their comments that helped improve the quality of the paper. This work is based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. This project acknowledges financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program HotMilk (grant agreement No. 865637), support from Bando per il Finanziamento della Ricerca Fondamentale 2022 dell’Istituto Nazionale di Astrofisica (INAF): GO Large program and from the Framework per l’Attrazione e il Rafforzamento delle Eccellenze (FARE) per la ricerca in Italia (R20L5S39T9). S.Mereghetti acknowledges financial support through the INAF grants “Magnetars” and “Toward Neutron Stars Unification”. SM acknowledges support from the grant NASA ADAP 80NSSC24K0666. CJ acknowledges the National Natural Science Foundation of China through grant 12473016, and the support by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550200). LS acknowledges funding from the grant entitled “Bando Ricerca Fondamentale INAF 2023”